OpenStudy (anonymous):

Which set of side lengths could form a triangle similar to a triangle with side lengths 11, 16, and 22? 22, 27, 33 22, 32, 42 33, 44, 66 33, 48, 66

jimthompson5910 (jim_thompson5910):

Let's assume choice A is the answer (it may or may not be)

jimthompson5910 (jim_thompson5910):

If choice A was the answer, then dividing the corresponding side lengths should give us the same number

jimthompson5910 (jim_thompson5910):

So 22/11 = 2 but 27/16 = 1.6875 so choice A is NOT a similar triangle with the given triangle

OpenStudy (anonymous):

is it D?

jimthompson5910 (jim_thompson5910):

it is D, nice job

hero (hero):

I was going to say Hint: Multiply the given triangle by 3

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)

jimthompson5910 (jim_thompson5910):

notice how 33/11 = 3 48/16 = 3 66/22 = 3 so we're tripling each given side to get the new corresponding side